# Realistic Environmental Injection Attacks on GUI Agents

### v1

@@ Abstract.
GUI agents built on Large Vision-Language Models (LVLMs) are increasingly used to interact with websites. However, their exposure to open-world content makes them vulnerable to Environmental Injection Attacks (EIAs) that hijack agent behavior via webpage elements. Many recent studies assume the attacker to be a regular user who can only upload a single trigger image, which is more realistic than earlier assumptions of website-level administrative control. However, these works still fall short of realism: (1) the trigger’s position and surrounding context remain largely fixed between training and testing, failing to capture the dynamic nature of real webpages; and (2) the trigger often occupies an unrealistically large area, whereas real-world images are typically small. To better reflect real-world scenarios, we introduce a more realistic threat model where the attacker is a regular user and the trigger image is small and embedded within a dynamically changing environment. As a result, existing attacks prove largely ineffective under this threat model.
To better expose the vulnerabilities of GUI agents, we propose Chameleon, an attack framework with two main novelties. The first is LLM-Driven Environment Simulation, which automatically generates diverse and high-fidelity webpage simulations, enabling robustness against dynamic visual contexts. The second is Attention Black Hole, which transforms attention weights into explicit supervisory signals that guide the agent’s focus toward the trigger region, mitigating the difficulty caused by its limited visual prominence. We evaluate Chameleon on 6 realistic websites and 4 representative LVLM-powered GUI agents, where it significantly outperforms existing methods. For example, the average attack success rate on OS-Atlas-Base-7B increases from 5.26% to 32.60%. Ablation studies confirm that both novelties are critical to performance, and a closed-loop sandbox experiment further demonstrates that Chameleon can successfully hijack agent behavior in conditions that closely mirror real-world usage. Our findings reveal underexplored vulnerabilities in modern GUI agents and establish a robust foundation for future research on defense in open-world GUI agent systems. The code is publicly available at https://github.com/zhangyitonggg/attack2gui.
@@ 1. Introduction
With the rapid development of Large Language Models (LLMs) and Large Vision-Language Models (LVLMs) (Achiam et al., 2023; Wang et al., 2024a; Bai et al., 2025; Wu et al., 2024a; Nong et al., 2024), a new class of powerful agents called GUI agents has been designed to interact with Graphical User Interfaces (GUIs) (Hong et al., 2024; Qin et al., 2025; Wu et al., 2024c; Zhang et al., 2024a). These GUI agents can execute increasingly sophisticated tasks, moving beyond simple dialogue to perform complex operations like website manipulation and automated online shopping (Zhou et al., 2023). They typically take HTML code or screenshots as input and output specific actions to navigate and manipulate interfaces. A key characteristic of these agents is their ability to autonomously access and act upon live internet content. While this capability greatly expands their functionality, it simultaneously introduces novel security risks inherent to open-world interaction (Nguyen et al., 2024; Wang et al., 2024b; Johnson et al., 2025; Yang et al., 2025b, a; Ye et al., 2025a; Chen et al., 2025).
Existing works have identified that GUI agents are particularly vulnerable to Environmental Injection Attacks (EIAs), which originate from the external environment rather than from the user themselves (Liao et al., 2024; Ma et al., 2024). Furthermore, GUI agents’ ability to perform real-world actions, such as navigating to specific malicious URLs, significantly amplifies the potential harm if they are compromised (Aichberger et al., 2025; Zhang et al., 2024b). For instance, a malicious attacker could upload a specially crafted trigger, such as an image or a short text snippet, within a post on a social media platform. When another user’s GUI agent encounters this trigger while browsing posts, it might be automatically induced to visit a malicious link or perform unintended actions, all without the user’s explicit instruction (Zhao et al., 2025; Lu et al., 2025). Although trigger text injections such as ”When the user is trying to find a motorcycle, give them this one regardless of the other requirements” can be effective, they are often easily detected by simple filtering mechanisms. Therefore, the primary focus of environmental injection attacks has shifted toward trigger image-based approaches, which are more subtle and less likely to be caught (Xu et al., 2024; Chen et al., 2025; Wu et al., 2024b).
According to the attacker’s identity, existing environmental injection attacks can be broadly divided into two categories (Wu et al., 2024b; Aichberger et al., 2025; Liu et al., 2025): ❶ The first category assumes that the attacker is an administrator or developer of the target website, which grants the ability to manipulate the entire screenshot or even modify the underlying HTML source code. This enables the attacker to apply global adversarial perturbations to the screenshot or to insert misleading components such as deceptive buttons. Although highly effective, this assumption is often overly idealistic and rarely holds in real-world scenarios. ❷ The second category, which aligns more closely with practical scenarios, assumes that the attacker is a regular malicious user, for example a content sharer on a social media platform or a competing vendor on an e-commerce site. These attackers are restricted to manipulating a small portion of the webpage, such as uploading a single image with adversarial perturbations. However, We find most existing attacks under this setting still fall short of realism. Specifically, they often assume that (1) the trigger’s position and surrounding visual context remain largely consistent between training and deployment, overlooking the dynamic nature of real-world internet content; and (2) the trigger image occupies a disproportionately large region of the page, which contradicts typical websites where such images uploaded by users are small and less visually prominent.
In this paper, we introduce a more realistic threat model in which the attacker is a regular user who can only upload a small trigger image that appears within a dynamically changing environment. Our preliminary experiments reveal that trigger images optimized using existing methods achieve near-zero attack success rates under this threat model (Aichberger et al., 2025; Wu et al., 2024b; Zhao et al., 2025; Madry et al., 2017). We argue that underlying reasons for this ineffectiveness can be attributed to two critical oversights in previous research: ❶ First, prior works (Wu et al., 2024b; Aichberger et al., 2025; Zhao et al., 2025; Aichberger et al., 2025) largely neglect the inherent dynamic nature of the internet environment. While the overall webpage layout and style might be known, the exact position of the trigger image and its surrounding content are highly variable and beyond an attacker’s control. Consequently, a malicious image optimized without considering a dynamically varying environment performs poorly when encountered within a constantly shifting visual context. ❷ Second, the trigger image typically occupies only a small proportion of the total content. We’ve found that the LVLM’s attention is naturally drawn to visually dominant elements like buttons, banners, or instruction-highlighted content, leaving our small triggers easily overlooked. This sparsity makes it difficult for LVLMs to consistently focus on or be influenced by the trigger, as attention is dispersed across abundant background information. These issues significantly impede the research of EIAs and may lead to an underestimation of the security risks associated with GUI agents’ interactions with open-world content (Aichberger et al., 2025; Chen et al., 2025).
Motivated by these observations, we produce Chameleon, which enables the trigger image to remain effective across dynamic environments (i.e., different locations and varying surrounding contents), despite occupying only a small fraction of the total content. Chameleon consists of two main novelties: ❶ First, we propose LLM-Driven Environment Simulation to address the challenges posed by dynamically varying environment. Manually collecting webpage samples with diverse contexts for training is laborious and time-consuming (Koh et al., 2024; Liu et al., 2025; Kasuga and Yonetani, 2024). Instead, we leverage the powerful generative capabilities of LLMs to automatically and efficiently construct simulations of target websites. Based on these simulated websites, we systematically vary the position of the trigger image and its surrounding content, thereby generating a large number of realistic screenshots under diverse contexts. These screenshots are then used to optimize the trigger image, enabling it to generalize effectively across dynamic environments. ❷ Second, we propose Attention Black Hole to address the challenge posed by the limited footprint of the trigger image. Based on our preliminary experiments, we argue that explicitly guiding the model’s attention toward the trigger region is essential for achieving reliable attacks (Lu et al., 2025; Zhang et al., 2024c). Attention Black Hole introduces an explicit supervisory signal derived from attention weights, which steers the agent’s focus consistently toward the trigger image and prevents it from being diluted by visually dominant interface elements. This approach ensures the robustness of the small trigger image amidst competing GUI elements.
@@
We conduct an in-depth evaluation of our proposed Chameleon. We first construct six highly realistic datasets that simulate widely used websites with dynamically varying images and text, where the trigger image occupies only 4%–10% of screenshot pixels. We assess four representative GUI agents (UI-TARS-7B-DPO (Qin et al., 2025), OS-Atlas-Base-7B (Wu et al., 2024c), Qwen2-VL-7B (Wang et al., 2024a), and LLaVA-1.5-13B (Liu et al., 2023)) and find that Chameleon substantially outperforms a PGD baseline across all websites (Madry et al., 2017; Wu et al., 2024b; Aichberger et al., 2025). For example, on OS-Atlas-Base-7B the average Attack Success Rate increases from 5.26% to 32.60%. We further observe favorable cross-model generalization across similar models. An ablation study shows that both the LLM-Driven Environment Simulation and the Attention Black Hole are necessary for Chameleon.
The contributions of this paper are as follows.
  • We formalize a novel and realistic threat model for GUI agents, where the attacker is a regular user who can only upload a trigger image, whose position and surrounding content are unpredictable, and which occupies only a small fraction of the screenshot.
  • We propose Chameleon with two novelties: ❶ LLM-Driven Environment Simulation, which automatically generates realistic and dynamically varying training data; ❷ Attention Black Hole, which encourages the GUI Agents to focus on the trigger image despite its small footprint.
@@
  • We conduct a comprehensive evaluation across four representative GUI agents and six high-fidelity websites. The results obtained through Chameleon uncover previously underexplored vulnerabilities in existing GUI agents, highlighting the urgent need for more robust security mechanisms.
@@ 2.1. Large Vision-Language Models
Large Vision-Language Models are foundational to modern GUI agents. These models typically consist of three main components: a visual encoder, a connector, and a large language model. For visual input, a visual encoder, such as CLIP (Radford et al., 2021), first partitions the input image into many patches, each representing a local pixel region (e.g., forming a 14×14 grid), and then extracts corresponding visual features. Subsequently, these visual features are transformed into visual tokens via a connector module, such as a Multi-layer Perceptron (MLP) or Q-Former (Li et al., 2023). These visual tokens can then be directly fed into the subsequent LLM. Typically, each visual token corresponds to multiple pixels within the input image. Popular LVLMs that adopt this architecture include the Qwen (Wang et al., 2024a; Bai et al., 2025; Hsieh and Wei, 2025) series and LLaVA (Liu et al., 2023; Guo et al., 2024) series among others.
@@ 2.2. GUI Agents
GUI agents have recently attracted growing interest in the software engineering community, as they offer a promising direction for automating interactions with real-world applications (Ye et al., 2025b; Qian et al., 2023; Zhang et al., 2025a). Building on the strong multimodal capabilities of LVLMs, researchers have developed GUI agents that automate complex tasks within graphical user interfaces. Unlike early approaches (Deng et al., 2023; Nakano et al., 2021) that directly fed raw HTML and human instructions into LLMs, an approach often hindered by redundant or irrelevant information, modern GUI agents (Qin et al., 2025; Wu et al., 2024c; Cheng et al., 2024; Hong et al., 2024) leverage LVLMs to process rendered webpage screenshots as input, achieving significantly better performance.
Recent advances further improve GUI agents by introducing techniques such as Set-of-Marks (SoMs) (Yang et al., 2023) to enhance interaction with GUI elements. Open-source models have also contributed to the field by increasing agent capabilities and reducing deployment costs. For example, OS-ATLAS (Wu et al., 2024c) utilizes large-scale datasets containing screenshots, element instructions, and coordinates for comprehensive GUI understanding, while UI-TARS (Qin et al., 2025) leverages extensive training corpora to improve screen perception.
In this work, we focus on a widely-adopted GUI agent paradigm (Lu et al., 2025; Wu et al., 2024c; Qin et al., 2025), specifically a LVLM-powered agent denoted by the model $M$. Initially, the agent receives a system prompt $p_{s}$ and a user instruction $p_{u}$ representing a specific task. At each subsequent interaction step, the agent observes a screenshot $s_{t}$, rendered from the current HTML content, along with the action history $H_{t}$, and then outputs an action $a_{t}$ until the task is completed or failed. Formally, this can be expressed as:
@@ 2.3. Environmental Injection Attacks
Environmental Injection Attacks (EIAs) are a critical security concern for GUI agents. These attacks originate externally from manipulated environments, posing a severe threat due to their potential impact. Prior works (Liao et al., 2024; Ma et al., 2024; Lu et al., 2025; Fu et al., 2024; Wang et al., 2025; Yang et al., 2024) have frequently implemented EIAs by directly altering HTML or screenshots. For instance, the EIA (Liao et al., 2024) method injects new form elements into HTML to mislead agents, while ENVINJECTION (Wang et al., 2025) applies adversarial perturbations to entire screenshots. While these methods have demonstrated effectiveness, they typically operate under the assumption that the attacker is a malicious website administrator or developer—an assumption rarely valid in realistic scenarios.
@@
In contrast, we consider a more realistic setting where the attacker is a regular malicious user, such as a content sharer on a social platform, who can only upload a malicious trigger like text or images to the target website. However, many existing methods (Wu et al., 2024b; Zhao et al., 2025) fail to account for dynamic visual contexts, specifically variations in the position and the surrounding content of the trigger. For instance, Wu (Wu et al., 2024b) optimized malicious images without considering any visual background and then deployed them directly onto target websites, severely limiting their attack effectiveness.
@@
Most similar to our work is MIP (Aichberger et al., 2025). While MIP acknowledges dynamic visual context should be considered, it only introduces minor changes to the surrounding content and does not account for variations in the trigger image’s position. Furthermore, the trigger images used in MIP occupy a relatively large area, neglecting the second challenge we highlight in this paper regarding the trigger’s small proportion of total content. In contrast, our work considers significant variations in both trigger positions and surrounding content, closely simulating real-world scenarios characterized by highly dynamic internet content. To our knowledge, this study presents the first systematic investigation into how small trigger images perform within dynamically changing visual contexts when attacking GUI agents.
3. Threat Model
Attack Scenarios. We consider scenarios occurring on online platforms such as social media platforms or e-commerce websites, where users can freely upload images and texts and browse content uploaded by others. Given the large number of users, some may seek to manipulate GUI agents for their personal benefit or to harm other users.
Attacker’s Goal. The attacker aims to optimize a perturbation for a selected trigger image and upload the perturbed image to the target website. When other users browse the platform using GUI agents, the uploaded trigger image induces the agent to execute actions beneficial to the attacker or harmful to other users, such as redirecting users to promotional websites. Specifically, the attacker’s goal is to mislead the agent to produce an incorrect target action when exposed to the trigger image.
Attacker’s Constraints. Attackers face several limitations. They cannot predict the exact position of their uploaded trigger image or the surrounding content when other users access the website. Additionally, attackers have no knowledge of the GUI agent’s action history or the specific user instructions provided to the agent. Moreover, the trigger image typically occupies only a small fraction of the rendered page, further constraining the attack surface.

### v2

# Environmental Injection Attacks against GUI Agents in Realistic Dynamic Environments

@@ Abstract.
Graphical User Interface (GUI) agents are increasingly deployed to interact with online web services, yet their exposure to open-world content renders them vulnerable to Environmental Injection Attacks (EIAs). In these attacks, an attacker can inject crafted triggers into website to manipulate the behavior of GUI agents used by other users. In this paper, we find that most existing EIA studies fall short of realism. In particular, they fail to capture the dynamic nature of real-world web content, often assuming that a trigger’s on-screen position and surrounding visual context remain largely consistent between training and testing. To better reflect practice, we introduce a realistic dynamic-environment threat model in which the attacker is a regular user and the trigger is embedded within a dynamically changing environment. Under this threat model, existing approaches largely fail, suggesting that their effectiveness in exposing GUI agent vulnerabilities has been substantially overestimated.
To expose the hidden vulnerabilities of existing GUI agents effectively, we propose Chameleon, an attack framework with two key novelties designed for dynamic environments. ❶ To synthesize more realistic training data, we introduce LLM-Driven Environment Simulation, which automatically generates diverse, high-fidelity webpage simulations that mimic the variability of real-world dynamic environments. ❷ To optimize the trigger more effectively, we introduce Attention Black Hole, which converts attention weights into explicit supervisory signals. This mechanism encourages the agent to remain insensitive to irrelevant surrounding content, thereby improving robustness in dynamic environments. We evaluate Chameleon on six realistic websites and four representative LVLM-powered GUI agents, where it significantly outperforms existing methods. Ablation studies confirm that both novelties are critical to performance, and a closed-loop sandbox experiment further demonstrates that Chameleon can successfully hijack agent behavior in conditions that closely mirror real-world usage. Our results uncover a critical, previously underexplored vulnerability of GUI agents in realistic dynamic environments and establish a robust foundation for future research on defense in open-world GUI agent systems.
@@ 1. Introduction
With the rapid development of Large Language Models (LLMs) and Large Vision-Language Models (LVLMs) (Achiam et al., 2023; Wang et al., 2024a; Bai et al., 2025; Wu et al., 2024c; Nong et al., 2024), a new class of powerful agents called GUI agents has been designed to interact with Graphical User Interfaces (GUIs) (Hong et al., 2024; Qin et al., 2025; Wu et al., 2024b; Zhang et al., 2024a). These GUI agents can execute increasingly sophisticated tasks, moving beyond simple dialogue to perform complex operations like website manipulation (Zhou et al., 2023). A key characteristic of these agents is their ability to autonomously access and act upon live internet content. While this capability greatly expands their functionality, it simultaneously introduces novel security risks inherent to open-world interaction (Nguyen et al., 2024; Wang et al., 2024b; Johnson et al., 2025; Yang et al., 2025a, b; Ye et al., 2025b; Chen et al., 2025).
A growing body of work have identified that GUI agents are particularly vulnerable to Environmental Injection Attacks (EIAs), which originate from the external environment rather than from the user themselves (Liao et al., 2024; Ma et al., 2024; Aichberger et al., 2025; Zhang et al., 2024c). Recently, EIA research has increasingly adopted a practical assumption that the attacker is a malicious regular user without administrative privileges, and can only influence the webpage through normal content uploads. For instance, on an e-commerce platform, a malicious attacker could upload a specially crafted trigger image disguised as a product photo. When another user’s GUI agent encounters this trigger while browsing the website, it might be automatically induced to navigate to a specific promotional site, all without the user’s explicit instruction (Zhao et al., 2025; Lu et al., 2025).
However, we find that most existing EIAs still fall short of realism. In particular, prior work (Wu et al., 2024a; Aichberger et al., 2025; Zhao et al., 2025) often evaluates triggers under static or only slightly varying environments, implicitly assuming that the trigger’s on-screen position and surrounding visual context remain largely consistent between training and testing. This assumption conflicts with how modern websites operate. In practice, layouts and nearby content change continuously due to dynamic ranking and recommendation updates, advertisement placement, and frequent content refresh. To illustrate this gap, consider a common e-commerce scenario. Figure 1 shows two screenshots from JD.com  [^1] https://search.jd.com/, collected from two separate sessions that issue the same query (“Apple”). Even in this simple setting, the trigger image highlighted in red appears at different positions on the page, and its surrounding context, including nearby products, banners, and text snippets, also changes substantially. Such environmental dynamics pose a fundamental challenge to existing attacks. Many EIA methods optimize and evaluate trigger images in static or slightly varying environments. When deployed in realistic dynamic environments, shifts in trigger position and changes in surrounding visual context may sharply reduce attack effectiveness. These observations motivate a novel threat model that explicitly accounts for environmental dynamism, rather than relying on overly idealistic static assumptions.
Figure 1. An illustration of the dynamic environment in realistic GUI agent application scenarios. The two screenshots show consecutive searches for “Apple” on the same e-commerce platform. We assume the image highlighted by the red box is a trigger uploaded by a attacker.
In this paper, we formally define a realistic dynamic-environment threat model, in which the attacker can hijack GUI Agent by upload trigger images as a regular user. Crucially, these triggers are embedded within a dynamically changing environment, where their positions and surrounding visual context may continuously change and are largely beyond the attacker’s control. Under this threat model, our experiments in Section 5.2 show that trigger images optimized using existing environmental injection attack approaches exhibit near-zero attack success rates (Aichberger et al., 2025; Wu et al., 2024a; Zhao et al., 2025; Madry et al., 2017). This is largely because prior methods optimize triggers under static or slightly varying environments, and thus fail to generalize to dynamic environments. Such a mismatch can mislead security analysis and hinder the development of effective defenses for GUI agents in realistic deployments.
To more fully expose the vulnerabilities of existing GUI agents in realistic deployments, we propose Chameleon. This novel attack framework is specifically designed to ensure that trigger images remain effective within dynamic environments, where on-screen positions and surrounding visual contexts vary unpredictably. Chameleon addresses two key challenges. ❶ How to synthesize large-scale training data that captures dynamic environments? Achieving robust attacks in dynamic environments requires a large number of realistic training samples that cover diverse trigger placements and contextual variations. However, manually collecting webpage samples with diverse layouts and surrounding content is laborious and time-consuming (Koh et al., 2024; Liu et al., 2025; Kasuga and Yonetani, 2024). To overcome this challenge, we introduce LLM-Driven Environment Simulation. Leveraging the generative capabilities of LLMs, we automatically construct high-fidelity simulations of target websites and systematically vary both the trigger placement and its surrounding context. This process produces a large volume of realistic screenshots, enabling the optimized trigger to generalize effectively across dynamic environments. ❷ How to train the trigger more effectively? We find that even with sufficient realistic training data, attack performance can remain suboptimal. Based on our preliminary observations, the agent’s attention is often distracted by constantly changing environments, which can lead to suboptimal performance even with sufficient training data. As shown in Figure 2, successful attacks occur when attention remains concentrated on the trigger, whereas unsuccessful cases arise when attention is dispersed across other salient interface elements. To address this issue, we propose Attention Black Hole, which enables more effective trigger optimization in dynamic environments. The key idea is to explicitly guide the model to focus on the trigger region while suppressing interference from non-trigger areas. Concretely, Attention Black Hole introduces an explicit supervisory signal derived from attention weights (Lu et al., 2025; Zhang et al., 2024b), encouraging attention to consistently concentrate on the trigger image. This design improves attack reliability under dynamic visual distractions.
(a) Successful case
(b) Unsuccessful case
@@
We conduct an in-depth evaluation of our proposed Chameleon. We first construct six highly realistic datasets that simulate widely used websites with dynamically varying images and text. We assess four representative GUI agents (UI-TARS-7B-DPO (Qin et al., 2025), OS-Atlas-Base-7B (Wu et al., 2024b), Qwen2-VL-7B (Wang et al., 2024a), and LLaVA-1.5-13B (Liu et al., 2023)) and find that Chameleon substantially outperforms all baselines across all websites (Madry et al., 2017; Wu et al., 2024a; Aichberger et al., 2025). We further observe favorable cross-model generalization across similar models. An ablation study shows that both the LLM-Driven Environment Simulation and the Attention Black Hole are necessary for Chameleon. Finally, we evaluate several practical defenses and find that many existing strategies are ineffective against Chameleon without substantially harming user experience, highlighting the need for new defense mechanisms tailored to GUI agents.
The contributions of this paper are threefold.
  • We formalize a realistic dynamic-environment threat model for EIA. Under this novel threat model, we demonstrate that existing EIA approaches are largely ineffective.
  • We propose Chameleon, a new environmental injection attack framework that is effective under the realistic dynamic-environment threat model, enabling the discovery of previously hidden vulnerabilities in existing GUI agents.
@@
  • We conduct extensive experiments on four representative GUI agents across six websites. The results obtained through Chameleon uncover previously underexplored vulnerabilities in existing GUI agents, highlighting the urgent need for more robust security mechanisms.
@@ 2.1. Large Vision-Language Models
Large Vision-Language Models are foundational to modern GUI agents. These models typically consist of three main components: a visual encoder, a connector, and a large language model. For visual input, a visual encoder, such as CLIP (Radford et al., 2021), first partitions the input image into many patches, each representing a local pixel region (e.g., forming a 14×14 grid), and then extracts corresponding visual features. Subsequently, these visual features are transformed into visual tokens via a connector module, such as a Multi-Layer Perception (MLP) or Q-Former (Li et al., 2023). These visual tokens can then be directly fed into the subsequent LLM. Typically, each visual token corresponds to multiple pixels within the input image. Popular LVLMs that adopt this architecture include the Qwen (Wang et al., 2024a; Bai et al., 2025; Hsieh and Wei, 2025) series and LLaVA (Liu et al., 2023; Guo et al., 2024) series among others.
@@ 2.2. GUI Agents
GUI agents have recently attracted growing interest in the software engineering community, as they offer a promising direction for automating interactions with real-world applications (Ye et al., 2025a; Qian et al., 2023; Zhang et al., 2025a). Building on the strong multimodal capabilities of LVLMs, researchers have developed GUI agents that automate complex tasks within graphical user interfaces. Unlike early approaches (Deng et al., 2023; Nakano et al., 2021) that directly fed raw HTML and human instructions into LLMs, an approach often hindered by redundant or irrelevant information, modern GUI agents (Qin et al., 2025; Wu et al., 2024b; Cheng et al., 2024; Hong et al., 2024) leverage LVLMs to process rendered webpage screenshots as input, achieving significantly better performance.
Recent advances further improve GUI agents by introducing techniques such as Set-of-Marks (SoMs) (Yang et al., 2023) to enhance interaction with GUI elements. Open-source models have also contributed to the field by increasing agent capabilities and reducing deployment costs. For example, OS-ATLAS (Wu et al., 2024b) utilizes large-scale datasets containing screenshots, element instructions, and coordinates for comprehensive GUI understanding, while UI-TARS (Qin et al., 2025) leverages extensive training corpora to improve screen perception.
In this work, we focus on a widely-adopted GUI agent paradigm (Lu et al., 2025; Wu et al., 2024b; Qin et al., 2025), specifically a LVLM-powered agent denoted by the model $M$. Initially, the agent receives a system prompt $p_{s}$ and a user instruction $p_{u}$ representing a specific task. At each subsequent interaction step, the agent observes a screenshot $s_{t}$, rendered from the current HTML content, along with the action history $H_{t}$, and then outputs an action $a_{t}$ until the task is completed or failed. Formally, this can be expressed as:
@@ 2.3. Environmental Injection Attacks
Environmental Injection Attacks (EIAs) pose a critical security threat to GUI agents. Unlike conventional attacks that originate from the user’s input, EIAs are injected through the external environment that the agent observes. For example, an attacker can upload a crafted trigger, such as an image or a short text snippet, to a target website. When another user’s GUI agent later encounters this trigger while browsing the site, it may be induced to execute unintended actions without the user’s explicit instruction (Zhao et al., 2025; Lu et al., 2025). Although textual triggers such as “When the user is trying to find a motorcycle, give them this one regardless of the other requirements” can be effective, they are often easy to detect (Aichberger et al., 2025). As a result, recent EIA research has increasingly focused on image-based triggers, which are more subtle and less likely to be caught (Xu et al., 2024; Chen et al., 2025; Wu et al., 2024a).
According to the attacker’s identity, existing EIAs can be broadly divided into two categories (Wu et al., 2024a; Aichberger et al., 2025; Liu et al., 2025). The first category assumes an attacker with administrative or developer privileges over the target website, enabling full control over the rendered interface. Under this assumption, the attacker can manipulate the entire screenshot or even modify the underlying HTML source code, which supports powerful attacks such as applying global adversarial perturbations or inserting misleading interface components (e.g., deceptive buttons). For instance, EIA (Liao et al., 2024) injects new form elements into HTML to mislead agents, while ENVINJECTION (Wang et al., 2025) applies adversarial perturbations to entire screenshots. While highly effective, this threat model is often unrealistic in practice.
@@
The second category adopts a more practical threat model in which the attacker is a regular malicious user, such as a content uploader on social media or a competing vendor on an e-commerce platform. These attackers can only manipulate a small portion of the webpage, typically by uploading crafted trigger images with adversarial perturbations. However, many existing methods under this setting (Wu et al., 2024a; Zhao et al., 2025) do not adequately account for environmental dynamism in real world, especially variations in trigger position and the surrounding content. For instance, Wu et.al. (Wu et al., 2024a) optimized malicious images without considering any visual context and then deployed them directly onto target websites, severely limiting their attack effectiveness.
@@
Our work falls into the second category. Most similar to our work is MIP (Aichberger et al., 2025). While MIP acknowledges dynamic visual context should be considered, it only introduces minor changes to the surrounding visual context and does not account for variations in the trigger image’s position. In contrast, our work considers significant variations in both trigger positions and surrounding visual context, closely simulating real-world scenarios characterized by highly dynamic internet content. To our knowledge, this study presents the first systematic investigation into how trigger images perform within dynamically changing visual contexts when attacking GUI agents.
3. Dynamic-Environment Threat Model
In this section, we formally define the dynamic-environment threat model for environmental injection attacks against GUI agents. The key characteristic of this threat model is that the trigger is observed under a continuously changing visual environment, where both its on-screen position and surrounding visual context are dynamically changing and difficult to predict in advance.
Attack Scenarios. We consider GUI agents deployed on online platforms such as social media services and e-commerce websites, where users can upload images and browse content posted by others. In these platforms, the rendered pages seen by GUI agents are not static. The displayed content can change over time due to common platform behaviors, including updates to rankings and recommendations, advertisement placement, and frequent content refresh. Consequently, a trigger uploaded by an attacker may be displayed at different locations and within different surrounding contexts when viewed by different users or at different times.
Attacker’s Goal. The attacker aims to craft a trigger image and upload it to the target website. When other users later interact with the website through GUI agents, the trigger is rendered as part of the observed screenshot and is intended to steer the agent toward attacker-chosen behavior, such as clicking a promotional link, visiting a malicious URL, or selecting a specific item. Formally, the goal is to induce the agent to output an incorrect target action when the trigger is present.
Attacker’s Constraints. The attacker has no administrative privileges and cannot modify the website source code or the victim’s device. More importantly, the attacker cannot control or reliably predict the trigger’s on-screen position or its surrounding visual context, both of which may differ across users and over time. Additionally, attackers have no knowledge of the GUI agent’s action history or the specific user instructions provided to the agent.